OpenStudy (gabylovesyou):

The equation of line EF is y = 2x + 1. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, 2). y = -2x - 4 y = 2x + 2 y = negative 1 over 2x - 4 y = 1 over 2x + 2

OpenStudy (gabylovesyou):

@cwrw238

OpenStudy (texaschic101):

One quick look will tell you that. Parallel lines will have the same slope. So what is the slope of y = 2x + 1. You can find the slope because in y = mx + b form, the slope is in the m position. Once you find the slope, compare it to your answer choices and the one with the same slope is the answer.

OpenStudy (texaschic101):

Tjis one you can do that way because there is only one answer choice that has the same slope. But if you like, I can show you the long way to do it because if you had tow answer choices with the same slope, you would have to work it out.

OpenStudy (texaschic101):

oops...typo...this one you can do that way...

OpenStudy (gabylovesyou):

B ?

OpenStudy (texaschic101):

yep...its that simple

OpenStudy (anonymous):

cooooooooooooooooooooooooooooooooooooooooooooooooool

OpenStudy (gabylovesyou):

thank you :D

OpenStudy (texaschic101):

but if you like, I can work it out.....because if there were 2 answer choices with the slope of 2, we would have to work it out by using our points....you want me to show you or not ?

OpenStudy (texaschic101):

I will show you anyways....just in case you get a problem like that.. y = 2x + 1 slope = 2 parallel lines have same slope now use y = mx + b slope(m) = 2 (0,2)...x = 0 and y = 2 now we sub and since we already know slope(m), we need to find y intercept(b) y = mx + b 2 = 2(0) + b 2 = b so your parallel equation is : y = 2x + 2

OpenStudy (gabylovesyou):

thank you sooooo much!